OpenStudy (anonymous):

The length of a rectangle is 2 foot less than 3 times its width. Find the dimensions of this rectangle if the area is 96 square feet.

OpenStudy (anonymous):

@nincompoop please help

OpenStudy (anonymous):

Ok...so start with the equation for area: LW= area You are given that L=3W-2 So substiture 3W-2 in for L you will get LW=96(area) (3W-2)W=96 expand 3W(W) -2W=96 3W^2-2W=96 3w^2-2W-96=0 solve.......

OpenStudy (anonymous):

I have gotten that far on my own, but I don't know what to do next. I don't understand how to factor it out. @aryandecoolest

OpenStudy (mokeira):

let width be w length= 3w-2 A=l*w 96=w(3w-2) \[96=3w ^{2}-2w\] \[3w ^{2}-2w-96=0\]

OpenStudy (anonymous):

Thank you, but I have gotten that far on my own I just don't know how to factor with the 3 infront of the w. Can you please help me more? @Mokeira

OpenStudy (anonymous):

ok...so here's what I got......(w-6)(3w+16), then set each term to zero and solve for w, these will be your solutions

OpenStudy (anonymous):

THANK YOU! I got w=6 and w=-16/3 @aryandecoolest

OpenStudy (princeharryyy):

16 ans

OpenStudy (anonymous):

Hope i helped a bit...!! yeah anytime )

OpenStudy (princeharryyy):

@klovic